OpenStudy (anonymous):

Ian solved the system of equations by making a substitution for y in the second equation. what equation did he get the result of that substitution? y= -3x+1 5x-2y=31

OpenStudy (anonymous):

So you notice the first equation y is already solved for you, all you need to do is plug in that expression into the second equation as follow: \[\huge 5x-2\color{red}{(-3x+1)}=31\]

OpenStudy (anonymous):

the answer choices are: A) 5x+6x+1=31 b) 5x+6x-2=31 c)-15x+5x=31 d)-15x-x=31

OpenStudy (anonymous):

would the answer be B?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what is the x coordinate of the solution to the systems of equations? 3x-2y=-30 6x+3y=3

OpenStudy (anonymous):

You can solve for one of the variables in one of the equations then do as we did above make a substitution and solve for x.

OpenStudy (anonymous):

i dont get it. i have a time limit on a quiz and almost of out of time.

OpenStudy (anonymous):

Sorry Jeremy but we do not help with quiz's here, that's for you to do on your own.

OpenStudy (anonymous):

Please do not ask further questions as it will require me to suspend you, thanks.